OpenStudy (anonymous):

(tomorrow exam so please give quick help) find the resultent (in components) and its magnitue p=[2,2,2] q=[-4,-4,0] u=[2,2,7] find v so that v,p,q,u are in equilibrium

rvc (rvc):

p=i+2j q=4j-k u=4i-3j v=6i+2j+4k

OpenStudy (anonymous):

@amistre64 @ganeshie8 @phi @myininaya @dumbcow @whpalmer4 @Mashy

OpenStudy (anonymous):

ok ?! thats all ?

rvc (rvc):

wat u want

rvc (rvc):

no dr

rvc (rvc):

magnitude=\[\sqrt{i^2+j^2+k^2}\]

OpenStudy (anonymous):

ok sry i will complete the question

rvc (rvc):

@BSwan

OpenStudy (anonymous):

she doesn't know what she wants xD

rvc (rvc):

1=sqrt5 rest u try this is magnitude

OpenStudy (dumbcow):

for equilibrium q = [0,0, -9] This way the Net force is zero in all directions